myininaya (myininaya):

ok polpak i got a question for you how to solve x^2=6x mod 8

myininaya (myininaya):

we can't do this x^2-6x-8n=0 for n integer why?

OpenStudy (anonymous):

ho ho ho. guess what you have to do?!

myininaya (myininaya):

i tried and the answer failed

OpenStudy (anonymous):

because 8 is composite. i will be quiet now

myininaya (myininaya):

you don't have to be satellite he was just showing me how to do one last night

OpenStudy (anonymous):

you got satellite all excited ;)

OpenStudy (anonymous):

you have to.... COMPLETE THE SQUARE, your favorite. that is why i got excited. because this is myininayas favorite job

myininaya (myininaya):

lol

myininaya (myininaya):

so thats why learning to complete the square over doing -b/2a is more important lol

OpenStudy (anonymous):

satelitte , how about middle of two roots? lol

myininaya (myininaya):

x^2-6x=0 mod 8 (x-3)^2=9 mod 8 (x-3)^2=1 mod 8

OpenStudy (anonymous):

really that is what you will do \[x^2-6x\equiv 0 (mod 8)\]

OpenStudy (anonymous):

there that took a while to find /equiv

OpenStudy (anonymous):

\[(x-3)^2+9\equiv 1(mod 8)\]

myininaya (myininaya):

\[\equiv \] :)

myininaya (myininaya):

adding nine twice on one side?

OpenStudy (anonymous):

added nine to both sides

OpenStudy (anonymous):

\[x^2-6x\equiv 0 (8)\] \[x^2-6x+9\equiv 9 (8)\] \[x^2-6x+9\equiv 1(8)\]

OpenStudy (anonymous):

so \[(x-3)^2\equiv 1 (8)\]

OpenStudy (anonymous):

which means you have several choices for \[x-3\] since any square is congruent to 1 mod 8 i believe

OpenStudy (anonymous):

well, that is provided it is odd.

OpenStudy (anonymous):

in any case you can just check now because you have only 7 numbers to check, but since x - 3 is odd it means x is even so you have a few choices

OpenStudy (anonymous):

you are good from here right? i love that you asked a question that required completing the square. made my day in fact!

myininaya (myininaya):

lol

OpenStudy (anonymous):

@polpak sorry if i butted in but i couldn't help myself

myininaya (myininaya):

ok thanks satellite :) for showing me how

myininaya (myininaya):

i can do the rest x=2,4,6,8

OpenStudy (anonymous):

I think 0 is a solution

myininaya (myininaya):

yep

myininaya (myininaya):

it is

myininaya (myininaya):

do all even numbers work?

OpenStudy (anonymous):

Actually I may be interpreting this differently.. Lemme see here..

myininaya (myininaya):

i think they do

OpenStudy (anonymous):

Was the question \(x^2 = (6x\ mod\ 8)\) or was it \(x^2 \equiv 6x (mod8)\)

myininaya (myininaya):

i believe lagrange meant the second one

OpenStudy (anonymous):

0 is a solution be because we really do not say 8 mod 8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then satellite's interpretation is the right one.

OpenStudy (anonymous):

if you are working mod 8 you are working with the numbers 0,1,2,3,4,5,6,7 or at least with their equivalence classes, depending on how pedantic you want to be

OpenStudy (anonymous):

I like to be very pedantic ;p

OpenStudy (anonymous):

ok then you are working with \[\bar{0},\bar{1},\bar{2},\bar{3}...\]

myininaya (myininaya):

\[(2n-3)^2 \equiv 1 \mod 8=> 4n^2-2n(2)(3)+9-1=8k\] \[=> 4n^2-12n+8 = 8k\] \[=>4(n^2-3n+2) = 8k ?\] k is integer by the way if we can show n^2-3n+2 is even then the answer is any even number

myininaya (myininaya):

n^2-3n+2=(n-1)(n-2) lets try induction for n=1 it is even for n=2 it is even for n=3 it is even ok now let's assume it is even for some integer k=> so we have for (k-1)(k-2)=2j (j is an integer) now we need to show it is even for k+1 (k+1-1)(k+1-2) =(k-1+1)(k-2+1) =(k-1)(k-2)+(k-1)1+(k-2)1 =2j + k-1+k-2 =2j+2k-3 =2(j+k)-3 but that means for k+1 it is odd so maybe not every even number

OpenStudy (anonymous):

what are you trying to show? i am lost

myininaya (myininaya):

that even even number is a solution to the congruence equation

myininaya (myininaya):

every even*

myininaya (myininaya):

i wanted to show n^2-3n+2 is even for any number if i could haven shown that then i could haven shown any even number is a solution to x^2=6x mod 8

myininaya (myininaya):

but my proof by induction failed so n^2-3n+2 is not even even for any number and therefore x^2=6x mod8 does not have solution x=2m for m is integer

myininaya (myininaya):

it only works for some even integers

myininaya (myininaya):

not all

OpenStudy (anonymous):

what hold on. you want to show \[n^2-3n+2\] is even if n is even?

myininaya (myininaya):

no i wanted to show n^2-3n+2 is even for any number

OpenStudy (anonymous):

oh if n is any number, not any even number

OpenStudy (anonymous):

ok got it. forget the 2

myininaya (myininaya):

if n^2-3n+2 is even for any number, then we could haven shown any even number is a solution to x^2=6x mod 8

OpenStudy (anonymous):

you are working too hard i think

OpenStudy (anonymous):

you can clearly ignore the 2 so you want to show that \[n^2-3n\] is even but \[n^2-3n=n(n-3)\] so if n is even you are done, and in n is odd then n-3 is even so you are done

OpenStudy (anonymous):

Number of solutions of P(x) = 0 (mod n) is number of integers b where P(b) = 0 (mod n) and b is a residue modulo n.

myininaya (myininaya):

\[(2n-3)^2 \equiv 1 \mod 8=> 4n^2-2n(2)(3)+9-1=8k \] \[=> 4n^2-12n+8 = 8k \] \[=>4(n^2-3n+2) = 8k ? \] this was what i working off

OpenStudy (anonymous):

anyway we already showed that \[n^2\equiv 6n (8)\] if n = 0, 2, 4, 6

OpenStudy (anonymous):

or maybe you were unhappy with the 'complete the square" solution and you are looking for another one

myininaya (myininaya):

no i wanted to find all the answers

myininaya (myininaya):

a general solution

OpenStudy (anonymous):

what do you mean by "all" you are working modulo 8 so there are 8 equivalence classes

OpenStudy (anonymous):

general solution is n = 0, 2, 4, 6, mod 8

myininaya (myininaya):

ok lol

OpenStudy (anonymous):

which is another way of saying n is even

myininaya (myininaya):

it doesn't work for all even numbers though i proved that it didn't

OpenStudy (anonymous):

name one

myininaya (myininaya):

i did a proof by induction that failed why do i have to name one

OpenStudy (anonymous):

because just because your proof failed doesn't mean it is not true. it means your proof was wrong. if you want to show it is false you don't show it is false by saying my proof didn't work. you show it is false by exhibiting a counter example.

OpenStudy (anonymous):

i mean if i cannot prove fermat's last theorem, it doesn't mean it is wrong!

myininaya (myininaya):

ok let me scroll up and look at what i did again

OpenStudy (anonymous):

btw these things are almost never proved by induction

myininaya (myininaya):

if the the answer is all even numbers then my proof should i have worked

myininaya (myininaya):

i see nothing wrong with my approach

OpenStudy (anonymous):

i like that argument. it is not a valid one, but i like it.

OpenStudy (anonymous):

i will prove now that any odd number is congruent to 1 modulo 8

myininaya (myininaya):

no i'm just saying not all even numbers are solution to x^2=x mod 8

myininaya (myininaya):

but there are a lot of even even numbers that are

myininaya (myininaya):

as we seen by pluggin in

OpenStudy (anonymous):

\[1^2=1=\equiv 1(8)\] \[3^2=9\equiv 1(8)\] \[5^2=25\equiv 1(8)\] \[7^2=49\equiv 1(8)\]that is my proof

OpenStudy (anonymous):

i thought it was \[n^2\equiv 6n (8)\]

myininaya (myininaya):

it is i forgot the 6

myininaya (myininaya):

my bad

OpenStudy (anonymous):

ok so what did we do? rewrite as \[(n-3)^2\equiv 1(8)\] and solve this

OpenStudy (anonymous):

which is the same thing as saying \[z^2\equiv 1(8)\] with n - 3 replaced by z

OpenStudy (anonymous):

we see from my list above that this means z = 1, 3, 5, 7

OpenStudy (anonymous):

which mean n =0, 2, 4, 6

OpenStudy (anonymous):

dont forget you are working mod 8, so your universe is {0,1,2,3,4,5,6,7}

OpenStudy (anonymous):

congruent mod 8 is an equivalence relation, so you are working with equivalence classes.

OpenStudy (anonymous):

like cosets. you have for example \[71 \in \bar{7}\]

myininaya (myininaya):

ok satellite give me another equation and let me see if i can solve it

OpenStudy (anonymous):

off the top of my head? ok

OpenStudy (anonymous):

\[x^2\equiv 2x (11)\]

OpenStudy (anonymous):

maybe there is no solution, i don't know.

myininaya (myininaya):

ok

OpenStudy (anonymous):

oh there are two, but easy ones. let me try again